Question 1 Determine the pH of a 0.75 M solution of CH5COOH. (Ka C6H5COOH) = 6.5 x 10-5).
Answer #1

[H+) ion concentration is given as, [++] = √kaXc Concentration of solutionc=0-75M [Ht] =14875x10-5 Ka= 6.5x10-5 [ht] = 10.75X6.5X10-5 [H+] = 6.988 10-3M pH = -log[ht] pH = -log (6.98x10-3) [pH = 2.16